At max speed an airplane travels 1680 miles on 5 hours. Flying with the wind it travels the same distance in 4 hours. What is the speed of the plane without the wind, and what is the wind speed.
Answer by TimothyLamb(4379)

x = speed of airplane in still air = 1680/5 = 336 mph
y = speed of wind
---
s = d/t
---
with the wind:
x + y = 1680/4
336 + y = 1680/4
y = 1680/4 - 336
y = 84
---
answer:
x = speed of airplane in still air 336 mph
y = speed of wind = 84 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php